OpenStudy (anonymous):

A rocket is launched from ground level with an initial vertical velocity (v0) of 176 ft/s. After how many seconds will the rocket hit the ground? (Hint: h(t) = -16t2 + v0t + h0).

OpenStudy (ash2326):

@tomtom777 You have the equation for the height of the rocket at any time t \[h(t)=-16t^2+v_0 t+h_0\] \[h_0= \text{height of the launch pad. Here it's 0, since rocket is launched from ground}\] \[v_0=\text{initial vertical velocity}\]

OpenStudy (ash2326):

Do you understand this?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Taking the initial height h0 =0 you get:\[h(t)=-16t^2+v_{0}t\]Now, since we are calling ground level h=0, we need to find t such that h(t)=0. We have:\[-16t^2+v_{0}t=0\]We are given the initial velocity:\[-16t^2+176t=0\]We can factor out a t and a factor of 16:\[16t(-t+11)=0\]We solve each factor separately. This equation is true is 16t=0 (i.e., t=0) and/or if -t+16=0 (i.e., t=16). Therefore, the rocket is at ground level at t=0 and at t=16 seconds :)

OpenStudy (anonymous):

oooops, should say -t+11=0, giving t=11 seconds. typo

OpenStudy (anonymous):

The rocket hits the ground after 11 seconds.